A firm faces the demand schedule q = 400 − 2p − p2. What price does it need
to charge to sell 100 units?
1
Expert's answer
2020-12-15T08:45:25-0500

"q = 100 \\

100 = 400-2p-p^2 \\

0 = 300-2p-p^2

-p^2-2p+300 = 0 \\

p_{1,2}=\frac{-(-2)±\sqrt{(-2)^2-4(-1) \times 300}}{2(-1)} \\

p_{1,2}= \frac{-(-2)± 2 \sqrt{301}}{2(-1)} \\

p_1 = -1 - \sqrt{301} = -18.35 \\

p_2 = \sqrt{301}-1= 16.35"

Answer: p = 16.35